A positive charge of 4.60 μC is fixed in place. From a distance of 3.00 cm a particle of mass 5.90 g and charge +3.00 μC is fired with an initial speed of 62.0 m/s directly toward the fixed charge. How close to the fixed charge does the particle get before it comes to rest and starts traveling away?
_____???

Please help me solve this ASAP and show work! Thanks!
Answer #1

Here,

let the closest distance is d

Using conservation of energy

0.50 * m * v^2 + k * q1 * q2/d1 = k * q1 * q2/d

0.50 * 0.0059 * 62^2 + 9 *10^9 * 4.6 *10^-6 * 3 *10^-6/0.03 = 9 *10^9 * 4.6 *10^-6 * 3 *10^-6/d

solving for d

d = 0.00802 m = 8.02 mm = 0.8 cm

the closest distance is 0.8 cm